OpenStudy (anonymous):

A circle has a radius of 8 inches. Find the area of a sector of the circle if the sector has an arc that measures 45°

mathslover (mathslover):

ok so first of all : what is the area of a sector :? do you know about it @ali1995

mathslover (mathslover):

so it is : \[\large{\frac{\theta}{2}r^2}\](in radians)

mathslover (mathslover):

but the value of angle is given in degrees so right that as : \[\large{\frac{\theta}{360}\pi r^2}\] ( in degrees)

mathslover (mathslover):

given radius = 8 inches and \(\theta = 45\)

mathslover (mathslover):

put the values : \[\LARGE{\frac{45}{360}\times \pi \times 8^2}\]

OpenStudy (anonymous):

ok thank you